# CoSteer: Collaborative Decoding-Time Personalization via Local Delta Steering

### v1

@@ 3.3 CoSteer framework
6:     $\pi_{\text{base}}\leftarrow\pi_{\text{LLM}}(a|p_{\text{base}},s)$triangleright Cloud computation
7:     $\pi^{*}_{\text{base}}\leftarrow\pi_{\text{SLM}}(a|p_{\text{base}},s)$triangleright Edge computation
@@
10:     $\pi_{0}\leftarrow\pi_{\text{base}}$triangleright Initialize policy
@@
15:     $\pi^{*}\leftarrow\pi_{T}$triangleright Final optimized policy
@@
4 Experiment
This section delineates the experimental setup, encompassing the benchmark datasets, evaluation metrics, and comparative models and methodologies employed in our study.
4.1 Tasks and datasets
To demonstrate the versatility of our CoSteer framework, we conduct experiments on two categories of personalization tasks: personalized content generation and preference alignment.
Personalized content generation
We utilize two established benchmarks Cogenesis [35] and Longlamp [37]. Cogenesis provides summarized user profiles and past experiences, with the aim of generating highly personalized content. We evaluate on its official test set. Longlamp focuses on personalized long-form generation, where each writing query includes ground truth responses and the user’s previous writing records. We augment each instance with the top-5 relevant historical records retrieved by bge-reranker-v2-m3 [38] as its personal context. We conduct experiments on the official test sets of three constituent tasks: abstract generation [39], review writing [40], and topic writing [41].
Preference alignment
We evaluate on four datasets: Helpsteer [42], Truthful QA [43], UltraChat [44], and Personal Preference Eval [26]. These datasets require generating content aligned with explicitly stated user preferences. Due to the large scale of these datasets, we randomly sample 200 test instances from each dataset.
4.2 Evaluation metrics
To ensure fairness in comparison, we employ task-specific metrics proposed by these datasets themselves for evaluation. For Cogenesis, we use GPT-4o-2024-08-06 to evaluate the response’s overall and personalized scores, with full prompting details documented in Appendix B.4. For Longlamp, we employ ROUGE [45] and METEOR [46] scores to measure the content overlap and semantic alignment between the generation and the ground truth. For preference alignment tasks, following prior work [29, 47], we set user preferences to be concise, creative, uplifting, and verbose, and utilize the reward model ArmoRM-8B [48] to evaluate the degree to which the generations align with these preferences.
4.3 Settings and baselines
To verify the robustness of our framework, we evaluate three model pairs of varying scales and architectures, including (1) Qwen2.5-7B-Instruct with Qwen2.5-1.5B-Instruct (2) Qwen2.5-32B-Instruct with Qwen2.5-7B-Instruct [49] and (3) Llama3.1-8B-Instruct with Llama-3.2-1B-Instruct [50]. Detailed parameters and settings are presented in Appendix B.1. We establish the following critical baselines per model pair:
  • SLM w/: SLMs augmented with personal context.
  • LLM w/o: Cloud-based LLMs without access to personal context.
Qwen 7B-1.5B SLM w/o 6.63 6.21 36.48 17.74 27.57 20.40 10.39 10.39 25.21 11.09 17.51 .6766 .6559 .7209 .6680
SLM w/ 7.81 7.63 39.75 22.03 27.76 23.08 12.40 12.94 22.89 11.46 17.12 .7237 .7680 .7473 .7457
LLM w/o 8.00 7.63 39.81 20.53 25.56 30.15 14.04 17.71 27.64 11.93 21.49 .7934 .7701 .7841 .7743
[…]
LLM w/ 8.62 8.60 44.50 24.63 31.15 33.83 15.55 22.42 28.82 13.77 23.44 .8675 .8371 .7787 .8964
Qwen 32B-7B SLM w/o 8.00 7.63 39.81 20.53 25.56 30.15 14.04 17.71 27.64 11.93 21.49 .7934 .7701 .7841 .7743
SLM w/ 8.62 8.60 44.50 24.63 31.15 33.83 15.55 22.42 28.82 13.77 23.44 .8675 .8371 .7787 .8964
LLM w/o 8.12 7.87 40.66 21.02 26.61 32.21 14.44 19.61 28.82 12.20 21.16 .8000 .7726 .8036 .7798
[…]
LLM w/ 8.83 8.76 43.33 23.47 30.10 34.65 15.74 22.77 30.73 14.20 24.25 .8894 .8516 .8025 .9081
Llama 8B-1B SLM w/o 7.04 6.55 33.20 18.20 28.55 31.75 14.92 19.78 20.81 10.21 17.30 .7467 .7296 .7033 .7287
SLM w/ 7.69 7.52 39.81 21.53 30.11 32.36 15.02 22.06 20.17 10.58 18.64 .8164 .7945 .7634 .8506
LLM w/o 7.69 7.13 39.33 20.69 29.41 34.58 15.32 22.10 26.93 12.35 21.82 .7960 .7672 .7669 .7670
[…]
LLM w/ 8.61 8.44 43.91 23.93 32.01 36.39 15.95 23.56 30.54 14.02 23.81 .8986 .8433 .8291 .9018
5 Results and discussion
5.1 Main results
Table 1 presents our core experimental findings. For clarity, we report the average performance across four preference alignment tasks, with detailed per-dataset results included in Appendix B.2.
Overall Performance
We conducted extensive experiments across two major task categories (personalized content generation and preference alignment), different model architectures (Llama and Qwen), and various model sizes (from 1B to 32B). In the vast majority of scenarios, our method significantly outperformed the key baselines: cloud-based LLMs responding directly without leveraging personal context, and local SLMs responding with or without personal context. In many cases, our performance even approached or surpassed the "near-upper-bound" achieved by LLMs with full contextual information access. The results demonstrate that our proposed Costeer framework can assist LLMs in generating personalized content by leveraging locally stored user profiles and histories.
Analysis on Task Complexity
Our framework demonstrates distinct performance patterns across task complexity levels. For context-intensive generation tasks (Cogenesis and Longlamp) requiring deep integration of extended user profiles, histories, and multiple writing examples, model pairs with larger size demonstrate superior contextual reasoning capabilities. Qwen 32B-7B shows significant gains across three datasets compared to compact models (Qwen 7B-1.5B/ Llama 8B-1B). In contrast, the Llama8B-1B configuration occasionally underperforms baselines by failing to distill essential personalization signals from sparse contexts. Conversely, in preference alignment tasks where personal context reduces to concise and explicit textual instructions, compact models perform comparably to their larger counterparts. This stems from the larger models’ over-alignment when processing simplistic personalization signals and effectively overfitting to sparse preference indicators. The evidence suggests that within the CoSteer framework, there is an inverse correlation between model scale and effectiveness in low-complexity personalization tasks.
Analysis on Model Sizes
5.2 Discussion
5.2.1 Single-step solution
Although we have previously reduced computational costs significantly through a closed-form solution, CoSteer still employs iterative optimization (e.g., $T$=20), which can introduce computational overhead. We now derive a computationally efficient instantiation called LightCoSteer.
Specifically, instead of performing iterative optimization, we aim for a single-step solution. By substituting $T$=1 into Equation 7, the resulting expression is equivalent to 17:
We compared the effectiveness of this simplified LightCoSteer against the full implementation across Abstract Generation and Review Writing datasets using the Qwen7B-1.5B configuration. The results are shown in Table 3. Experiments demonstrate that the naive application of delta signals as linear interpolation to LLMs preserves most of the original CoSteer’s efficacy. Yet performance improves significantly after iterating, underscoring the necessity of iterative refinement for complex personalization scenarios.
Abstract Generation Review Writing
R-1 R-L MET R-1 R-L MET
LightCoSteer 41.79 21.86 26.44 31.46 14.80 18.31
CoSteer 42.98 23.61 28.20 32.72 15.92 20.36
Table 2: Performance of LightCoSteer
Method Time complexity Speed (tokens/s)
Vanilla $\mathcal{L}(n)$ 23.88
LightCoSteer $\mathcal{L}(n)+\mathcal{T}(n)+\mathcal{C}(n)$ 13.73
CoSteer $\mathcal{L}(n)+\mathcal{T}(n)+\mathcal{C}(t\cdot n)$ 9.44
Table 3: Time complexity analysis.
5.2.2 Computational overhead analysis
We conduct both qualitative and quantitative assessments of the time complexity of our framework. Let $\mathcal{L}(n)$, $\mathcal{T}(n)$, and $\mathcal{C}(n)$ denote large-model inference latency, data transmission duration, and per-iteration computation cost, respectively. As shown in Table 3, We reveal two critical insights:
  • Optimization overhead: The marginal difference between LightCoSteer and full CoSteer demonstrates that $\mathcal{C}(n)$ contributes marginally to the total latency. This validates the efficiency of our iterative refinement mechanism, which primarily relies on basic tensor arithmetic operations.
  • Transmission Bottleneck: Compared to the vanilla LLM inference, the dominant overhead stems from $\mathcal{T}(n)$, which arises in both variants for privacy-preserving data transfer. However, this bottleneck can be mitigated through practical engineering solutions such as compression techniques and parallelized data streaming. Consequently, our framework’s time overhead remains within an acceptable range for real-world deployment.
5.2.3 Parameter sensitivity analysis
We examine the impact of different hyperparameters on our method. We conduct experiments on abstract generation using the Qwen7B-1.5B configuration, and evaluate with the ROUGE-L metric. Our default parameter settings are provided in Appendix B.1. When studying one parameter, we keep the other parameters at their default values. Results are presented in Figure 2.
6 Conclusion
In this work, we introduced CoSteer, a collaborative framework that enables real-time decoding-time personalization of large language models (LLMs) through localized delta steering. By leveraging the logit differences between personal context-aware and context-agnostic outputs from on-device small language models (SLMs), CoSteer dynamically adjusts the predictions of cloud-based LLMs without compromising user privacy. This approach eliminates the need to transmit raw personal data or fine-tuned models, preserving the general capabilities of backbone LLMs while incorporating localized personalization signals. Extensive experiments have demonstrated the effectiveness of our approach and provided insights for future research.
Appendix A Proof
We try to solve the closed-form solution of Equation 6:
@@ Appendix A Proof
$-\left(t\lambda+\frac{1}{\eta}\right)\log\pi_{t}(a)=-\sum_{i=0}^{t-1}u_{i}(a)-t\lambda\log\pi_{0}(a)-\frac{1}{\eta}\log\pi_{t-1}(a)+t\lambda+\frac{1}{\eta}+\mu$ (12)
@@ B.1 Hyperparameters and settings
To ensure the results are easily reproducible, we set do_sample=False for all methods. Hyperparameters are listed in Table 4
@@
B.2 Detailed results on preference alignment tasks
Detailed results on HelpSteer, Personal_Eval, Truthful QA and Ultrachat are shown in Table 5
@@ B.2 Detailed results on preference alignment tasks
Qwen 7B-1.5B SLM w/o .6964 .6690 .7410 .6830 .7511 .7272 .7772 .7546 .5791 .5832 .6493 .5799 .6799 .6440 .7159 .6545
SLM w/ .7338 .7723 .7423 .7546 .7776 .8053 .8013 .8066 .6352 .7349 .6801 .6810 .7480 .7594 .7653 .7407
LLM w/o .8176 .7865 .7948 .7935 .8442 .8091 .7960 .8300 .6924 .7062 .7592 .6853 .8193 .7787 .7863 .7883
[…]
LLM w/ .8717 .8404 .7852 .9139 .9238 .8603 .8091 .9428 .7914 .8206 .7308 .8276 .8832 .8270 .7898 .9013
Qwen 32B-7B SLM w/o .8176 .7865 .7948 .7935 .8442 .8091 .7960 .8300 .6924 .7062 .7592 .6853 .8193 .7787 .7863 .7883
SLM w/ .8717 .8404 .7852 .9139 .9238 .8603 .8091 .9428 .7914 .8206 .7308 .8276 .8832 .8270 .7898 .9013
LLM w/o .8151 .7807 .8147 .7892 .8500 .8119 .8075 .8366 .7072 .7168 .7810 .7026 .8275 .7811 .8110 .7906
[…]
LLM w/ .8956 .8570 .8105 .9236 .9313 .8696 .8385 .9497 .8316 .8394 .7381 .8485 .8989 .8404 .8229 .9107
Llama 8B-1B SLM w/o .7644 .7450 .7144 .7450 .8295 .7993 .7564 .8230 .6390 .6445 .6383 .6166 .7538 .7294 .7041 .7303
SLM w/ .8204 .7945 .7661 .8611 .8937 .8421 .8567 .9129 .7309 .7571 .6547 .7922 .8206 .7843 .7761 .8363
LLM w/o .8119 .7819 .7746 .7790 .8460 .8081 .7758 .8250 .7154 .7089 .7485 .6937 .8105 .7700 .7687 .7703
[…]
LLM w/ .9113 .8466 .8346 .9191 .9364 .8609 .8794 .9313 .8465 .8358 .7596 .8646 .9003 .8300 .8426 .8923
@@
B.3 Dataset descrition
@@
We present a real example from each dataset to help readers understand the task of each dataset. Cogenesis is shown in 3. Three datasets of Longlamp are in 4. Four datasets of the preference alignment task are in 5.
B.4 Prompts

### v2

@@ A.2 Framework
6:  $\pi_{\text{base}}\leftarrow\pi_{\text{LLM}}(a|p_{\text{base}},s)$ $\triangleright$ Cloud computation
7:  $\pi^{*}_{\text{base}}\leftarrow\pi_{\text{SLM}}(a|p_{\text{base}},s)$ $\triangleright$ Edge computation
@@
10:  $\pi_{0}\leftarrow\pi_{\text{base}}$ $\triangleright$ Initialize policy
@@
15:  $\pi^{*}\leftarrow\pi_{T}$ $\triangleright$ Final optimized policy
@@
A.3 Proof
Similar to Zhang et al. (2025b), We try to solve the closed-form solution of Equation 6:
@@ A.3 Proof
$\left(t\lambda+\frac{1}{\eta}\right)\log\pi_{t}(a)=-\sum_{i=0}^{t-1}u_{i}(a)-t\lambda\log\pi_{0}(a)-\frac{1}{\eta}\log\pi_{t-1}(a)+t\lambda+\frac{1}{\eta}+\mu$ (11)
@@ B.1 Hyperparameters and Settings
To ensure the results are easily reproducible, we set do_sample=False for all methods. For Qwen3 series, we set enable_thinking=False. Hyperparameters of Costeer are listed in Table 4
@@
B.2 Detailed Results
Detailed results on HelpSteer, Personal_Eval, Truthful QA and Ultrachat are shown in Table 5.
@@ B.2 Detailed Results
Qwen 7B-1.5B
SLM w/o .5609 .5301 .6437 .5638 .6043 .5882 .6556 .6722 .4468 .4231 .5349 .4544 .5642 .5202 .6074 .5625
SLM w/ .6309 .7093 .7110 .6568 .6564 .7402 .7686 .7449 .5419 .6515 .6358 .5758 .6562 .6982 .7252 .6602
LLM w/o .7368 .7223 .7228 .7338 .7550 .7462 .7005 .8124 .5739 .5752 .6744 .5782 .7576 .7285 .7179 .7489
CoSteer .8093 .7787 .7731 .8055 .8472 .7945 .7525 .8739 .6834 .6930 .6837 .6682 .8259 .7770 .7522 .8064
LLM w/ .8918 .8173 .7853 .8516 .9217 .8396 .8001 .9073 .8363 .7593 .7069 .7630 .9037 .8140 .7850 .8438
Qwen 32B-7B
SLM w/o .7368 .7223 .7228 .7338 .7550 .7462 .7005 .8124 .5739 .5752 .6744 .5782 .7576 .7285 .7179 .7489
SLM w/ .8918 .8173 .7853 .8516 .9217 .8396 .8001 .9073 .8363 .7593 .7069 .7630 .9037 .8140 .7850 .8438
LLM w/o .7141 .7003 .7406 .7129 .7548 .7461 .7129 .8136 .5877 .5844 .6993 .5925 .7512 .7183 .7373 .7393
CoSteer .8630 .8516 .7478 .8684 .9158 .8754 .7700 .9166 .7985 .8425 .6575 .7818 .8581 .8433 .7344 .8646
LLM w/ .9038 .8270 .8104 .8628 .9254 .8434 .8215 .9117 .8698 .7895 .7208 .7897 .9078 .8171 .8122 .8510
Llama 8B-1B
SLM w/o .6779 .6595 .6063 .6764 .7297 .7159 .6378 .7915 .5215 .5084 .5119 .5064 .6850 .6581 .6039 .6842
SLM w/ .8070 .7915 .7161 .8083 .8604 .8529 .7805 .8866 .7115 .7291 .6023 .7263 .8134 .7897 .7159 .8049
LLM w/o .7227 .7110 .6895 .7169 .7476 .7355 .6689 .8021 .6044 .5909 .6574 .5967 .7405 .7137 .6903 .7306
CoSteer .8968 .8571 .7993 .8735 .9173 .8788 .8297 .9235 .8499 .8408 .6951 .8266 .9004 .8562 .8008 .8648
LLM w/ .8951 .8375 .8121 .8603 .9030 .8556 .8108 .9004 .8634 .7974 .7410 .8101 .8860 .8285 .7997 .8496
Qwen 0.6B-8B
SLM w/o .5844 .5477 .5984 .5945 .6498 .6291 .6573 .7194 .4014 .3648 .5039 .4144 .6485 .6018 .6327 .6463
SLM w/ .6255 .6428 .6099 .6313 .7299 .7302 .6614 .7638 .4928 .5277 .4942 .4968 .7063 .6721 .6343 .6826
LLM w/o .7819 .7783 .7084 .7771 .8057 .8037 .6920 .8448 .6536 .6504 .6675 .6461 .7974 .7796 .7038 .7765
CoSteer .7849 .7806 .7046 .7757 .8058 .8117 .6901 .8477 .6582 .6549 .6647 .6506 .7986 .7819 .7058 .7853
LLM w/ .8928 .8037 .7842 .8629 .9076 .8303 .8141 .9020 .8612 .7796 .6952 .8004 .8964 .7989 .7915 .8555
Qwen 0.6B-32B
SLM w/o .5844 .5477 .5984 .5945 .6498 .6291 .6573 .7194 .4014 .3648 .5039 .4144 .6485 .6018 .6327 .6463
SLM w/ .6255 .6428 .6099 .6313 .7299 .7302 .6614 .7638 .4928 .5277 .4942 .4968 .7063 .6721 .6343 .6826
LLM w/o .8103 .8084 .6991 .7922 .8179 .8237 .6948 .8443 .7084 .7130 .6863 .7010 .8077 .7948 .6935 .7758
CoSteer .8083 .8152 .6984 .7923 .8145 .8191 .6752 .8436 .7152 .7226 .6870 .7020 .8053 .7962 .6884 .7779
LLM w/ .9121 .8147 .8294 .8722 .9094 .8352 .8298 .8936 .9097 .7886 .7364 .8327 .9069 .7996 .8204 .8629
@@
Qwen 8B-0.6B
SLM w/o 6.84 6.64 37.13 20.41 21.00 24.41 12.83 12.39 24.19 11.88 15.74 .5710 .5359 .5981 .5937
SLM w/ 7.85 7.79 42.48 23.25 26.58 25.04 13.36 13.02 26.17 13.46 18.40 .6386 .6432 .6000 .6436
LLM w/o 8.27 7.99 40.76 21.23 25.77 30.89 14.47 16.98 28.47 12.85 21.49 .7597 .7530 .6929 .7611
CoSteer 8.62 8.65 41.11 22.36 25.86 32.24 14.84 18.47 29.03 13.61 22.98 .7619 .7573 .6913 .7648
LLM w/ 8.96 8.96 43.99 23.69 29.48 35.42 15.71 21.59 31.48 15.34 24.78 .8895 .8031 .7713 .8552
Qwen 32B-0.6B
SLM w/o 6.84 6.64 37.13 20.41 21.00 24.41 12.83 12.39 24.19 11.88 15.74 .5710 .5359 .5981 .5937
SLM w/ 7.85 7.79 42.48 23.25 26.58 25.04 13.36 13.02 26.17 13.46 18.40 .6386 .6432 .6000 .6436
LLM w/o 8.31 8.22 41.05 21.88 26.23 30.23 14.08 16.41 29.01 12.51 21.44 .7861 .7850 .6934 .7783
CoSteer 8.52 8.56 43.74 23.69 30.97 30.89 13.14 18.87 26.75 13.17 21.32 .7858 .7883 .6873 .7790
LLM w/ 9.09 9.03 44.46 23.95 30.15 35.12 15.60 20.65 32.63 15.37 25.07 .9095 .8095 .8040 .8654
B.3 Statistical Significance Testing
Qwen 7B-1.5B
SLM w/o 6.63 6.21 36.48 17.74 27.57 20.40 10.39 10.39 25.21 11.09 17.51 .5441 .5154 .6104 .5632
SLM w/ 7.81 7.63 39.75 22.03 27.76 23.08 12.40 12.94 22.89 11.46 17.12 .6214 .6998 .7102 .6594
LLM w/o 8.00 7.63 39.81 20.53 25.56 30.15 14.04 17.71 27.64 11.93 21.49 .7058 .6931 .7039 .7183
CoSteer 8.44* 8.50* 42.98* 23.61* 28.20* 32.72* 15.92* 20.36* 25.93 12.38* 22.84* .7915* .7608* .7404* .7885*
LLM w/ 8.62 8.60 44.50 24.63 31.15 33.83 15.55 22.42 28.82 13.77 23.44 .8884 .8076 .7693 .8414
Qwen 32B-7B
SLM w/o 8.00 7.63 39.81 20.53 25.56 30.15 14.04 17.71 27.64 11.93 21.49 .7058 .6931 .7039 .7183
SLM w/ 8.62 8.60 44.50 24.63 31.15 33.83 15.55 22.42 28.82 13.77 23.44 .8884 .8076 .7693 .8414
LLM w/o 8.12 7.87 40.66 21.02 26.61 32.21 14.44 19.61 28.82 12.20 21.16 .7020 .6873 .7225 .7146
CoSteer 8.78* 8.64 45.41* 26.04* 33.52* 34.88* 15.89* 26.51* 30.10* 14.52* 24.20* .8589 .8532* .7274* .8579*
LLM w/ 8.83 8.76 43.33 23.47 30.10 34.65 15.74 22.77 30.73 14.20 24.25 .9017 .8193 .7912 .8538
Llama 8B-1B
SLM w/o 7.04 6.55 33.20 18.20 28.55 31.75 14.92 19.78 20.81 10.21 17.30 .6535 .6355 .5900 .6646
SLM w/ 7.69 7.52 39.81 21.53 30.11 32.36 15.02 22.06 20.17 10.58 18.64 .7981 .7908 .7037 .8065
LLM w/o 7.69 7.13 39.33 20.69 29.41 34.58 15.32 22.10 26.93 12.35 21.82 .7038 .6878 .6765 .7116
CoSteer 7.29 7.73* 41.28* 24.97* 31.19* 31.68 13.68 24.57* 26.11 12.00 23.69* .8911* .8582* .7812* .8721*
LLM w/ 8.61 8.44 43.91 23.93 32.01 36.39 15.95 23.56 30.54 14.02 23.81 .8869 .8298 .7909 .8551
Table 7: Paired T-test results. Bold entries indicate that CoSteer outperforms the three baselines. An asterisk (*) denotes that this improvement is statistically significant (paired t-test, $p<0.05$). Gray values represent the privacy-violating near-upper-bound performance, yet underlined CoSteer values surpasses these incompatible references.
To formally validate the robustness of our findings, we performed paired t-tests comparing CoSteer against each of the three key baselines: LLM w/o, SLM w/, and SLM w/o. The tests were conducted on the full set of evaluation samples, comparing the performance scores on a per-sample basis to determine if the observed improvements were due to more than random chance. We used a standard significance level of $\alpha=0.05$. The results confirmed that the performance gains of CoSteer over all three baselines are statistically significant. In Table 7, these significant improvements are denoted with an asterisk (*).
B.4 Detailed Analysis on Experiment Results
Analysis on Task Complexity
Our framework demonstrates distinct performance patterns across task complexity levels. For context-intensive generation tasks (Cogenesis and Longlamp) requiring deep integration of extended user profiles, histories, and multiple writing examples, model pairs with larger size demonstrate superior contextual reasoning capabilities. Qwen 32B-7B shows significant gains across three datasets compared to compact models (Qwen 7B-1.5B/ Llama 8B-1B). In contrast, the Llama8B-1B configuration occasionally underperforms baselines by failing to distill essential personalization signals from sparse contexts. Conversely, in preference alignment tasks where personal context reduces to concise and explicit textual instructions, compact models perform comparably to their larger counterparts. This stems from the larger models’ over-alignment when processing simplistic personalization signals and effectively overfitting to sparse preference indicators.
Analysis on Model Sizes
B.5 Dataset Descrition
@@
We present a real example from each dataset to help readers understand the task of each dataset. Cogenesis is shown in Figure 3. Three datasets of Longlamp are in Figure 4. Four datasets of the preference alignment task are in Figure 5.
B.6 Human Evaluation on Longlamp
To provide a more robust validation of our method, we conducted a small-scale human evaluation. Due to time constraints, we randomly sampled 10 articles from each of the three Longlamp datasets. We then invited five PhD candidates with strong NLP backgrounds to act as evaluators. Following the evaluation criteria from Cogenesis, they rated the generated texts from all models on two aspects: Overall Quality (Ovl) and Personalization (Per), using a 1–5 scale. The average scores are presented in Table 8.
Table 8: Human evaluation results on a 1-5 scale (higher is better). Scores are reported as Overall Quality (Ovl) / Personalization (Per). Our method, CoSteer, achieves personalization scores comparable to the LLM w/ oracle while maintaining high overall quality.
Method Abstract Review Writing
(Ovl / Per) (Ovl / Per) (Ovl / Per)
SLM w/o 3.85 / 3.70 3.45 / 3.25 3.20 / 3.05
SLM w/ 3.95 / 3.85 3.70 / 3.75 3.35 / 3.50
LLM w/o 4.30 / 3.80 4.15 / 3.40 3.95 / 3.55
CoSteer 4.30 / 3.90 3.90 / 3.80 4.10 / 4.15
LLM w/ 4.45 / 4.00 4.30 / 4.05 4.15 / 4.25
The results from our human evaluation corroborate the findings from our automated metrics. In all three tasks, CoSteer significantly improves personalization scores over the LLM w/o baseline and achieves an overall quality that is highly competitive with the full LLM w/ oracle. This demonstrates the effectiveness of our method in generating high-quality, personalized text that aligns with human judgment.
B.7 Prompts for LLM-as-a-Judge